# Incorporating Arbitrary Matrix Group Equivariance into KANs

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have seen great success in scientific domains thanks to spline activation functions, becoming an alternative to Multi-Layer Perceptrons (MLPs). However, spline functions may not respect symmetry in tasks, which is crucial prior knowledge in machine learning. Previously, equivariant networks embed symmetry into their architectures, achieving better performance in specific applications. Among these, Equivariant Multi-Layer Perceptrons (EMLP) introduce arbitrary matrix group equivariance into MLPs, providing a general framework for constructing equivariant networks layer by layer. In this paper, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), a method for incorporating matrix group equivariance into KANs, aiming to broaden their applicability to more fields. First, we construct gated spline basis functions, which form the EKAN layer together with equivariant linear weights. We then define a lift layer to align the input space of EKAN with the feature space of the dataset, thereby building the entire EKAN architecture. Compared with baseline models, EKAN achieves higher accuracy with smaller datasets or fewer parameters on symmetry-related tasks, such as particle scattering and the three-body problem, often reducing test MSE by several orders of magnitude. Even in non-symbolic formula scenarios, such as top quark tagging with three jet constituents, EKAN achieves comparable results with EMLP using only $26\%$ of the parameters, while KANs do not outperform MLPs as expected.
@@ 1 Introduction
Kolmogorov-Arnold Networks (KANs) [37, 36] are a novel type of neural network inspired by the Kolmogorov-Arnold representation theorem [46, 5], which offers an alternative to Multi-Layer Perceptrons (MLPs) [26, 11, 28]. Unlike MLPs, which utilize fixed activation functions on nodes, KANs employ learnable activation functions on edges, replacing the linear weight parameters entirely with univariate functions parameterized as splines [14]. On the other hand, each layer of KANs can be viewed as spline basis functions followed by a linear layer [16]. This architecture allows KANs to achieve better accuracy in symbolic formula representation tasks compared with MLPs, particularly in function fitting and scientific applications. Subsequent works based on KANs have demonstrated superior performance in other areas, such as sequential data [47, 20, 19, 55], graph data [6, 15, 31, 58], image data [8, 7, 1, 34, 41, 3], and so on.
@@
However, KANs themselves perform poorly on non-symbolic formula representation tasks [56]. One reason for this is that splines struggle to respect data type and symmetry, both of which play important roles in machine learning. Many recent works utilize symmetry in data to design network architectures, achieving better efficiency and generalization on specific tasks. For example, Convolutional Neural Networks (CNNs) [32] and Group equivariant Convolutional Neural Networks (GCNNs) [9] leverage translational and rotational symmetries in image data, while DeepSets [57] and equivariant graph networks [38, 30, 40] exploit the permutation symmetry in set and graph data.
Equivariant Multi-Layer Perceptrons (EMLP) [17] propose a general method that allows MLPs to be equivariant with respect to arbitrary matrix groups for specific data types, thereby unifying the aforementioned specialized network architectures. They construct the equivariance constraint using discrete and infinitesimal generators of the matrix group, and then obtain equivariant linear weights by solving this constraint. Next, they apply gated nonlinearities [51], which use the additional output of the previous layer as gate scalars. Due to the generality of matrix groups, EMLP outperforms non-equivariant baselines on several datasets with complex symmetries.
@@
Inspired by EMLP, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), which embed matrix group equivariance into KANs. By specifying the data type and symmetry, EKAN can serve as a general framework for applying KANs to various areas. In Sec. 2, we introduce the preliminary knowledge of group theory. In Sec. 3, we summarize related works. In Sec. 4, we construct a layer of EKAN. We add gate scalars to the input and output space of each layer, and define gated spline basis functions between the input and post-activation space. To ensure equivariance when linearly combining gated basis functions, we construct the equivariant constraint and solve for the equivariant linear weights similar to EMLP. In Sec. 5, we build the entire EKAN architecture. We insert a lift layer before the first layer and discard the gate scalars from the output of the final layer, so that the input and output space of EKAN can be consistent with the original dataset. In Sec. 6, we evaluate EKAN on tasks with known symmetries. We show that EKAN can achieve higher accuracy than baseline models with smaller datasets or fewer parameters. In Sec. 7, we conclude this work. In Fig. 1, we compare the architectures of MLPs, EMLP, KANs, and EKAN.
@@
Figure 1: Comparison of the architectures of Multi-Layer Perceptrons (MLPs), Equivariant Multi-Layer Perceptrons (EMLP), Kolmogorov-Arnold Networks (KANs), and Equivariant Kolmogorov-Arnold Networks (EKAN).
In summary, our contributions are as follows:
  • We propose EKAN, an architecture that makes KANs equivariant to matrix groups. To our knowledge, EKAN is the first attempt to combine equivariance with KANs, and we expect that it can serve as a general framework to broaden the applicability of KANs to more areas.
  • We specify the space structures of the EKAN Layer and define gated spline basis functions. We theoretically prove that gated basis functions can ensure equivariance between the gated input space and the post-activation space. Then, we insert a lift layer to preprocess the raw input feature, which aligns the input space of EKAN with the feature space of the dataset.
@@
  • Experiments on tasks with matrix group equivariance, such as particle scattering and the three-body problem, demonstrate that EKAN often significantly outperforms baseline models, even with smaller datasets or fewer parameters. In the task of non-symbolic formula representation, where KANs are not proficient, such as top quark tagging with three jet constituents, EKAN can still achieve comparable results with EMLP while using only $26\%$ of the parameters.
@@ 2.1 Groups and generators
where $\{A_{i}\}_{i=1}^{D}$ are called infinitesimal generators and $\{h_{i}\}_{i=1}^{M}$ are called discrete generators. We introduce common matrix groups and their generators in Supplementary Material.
2.2 Group representations
The group representation $\rho_{V}:\widetilde{G}\rightarrow\mathrm{GL}(m)$ maps group elements to $m\times m$ invertible matrices, which describes how group elements act on the vector space $V=\mathbb{R}^{m}$ through linear transformations. For $g_{1},g_{2}\in\widetilde{G}$ it satisfies $\rho_{V}(g_{1}g_{2})=\rho_{V}(g_{1})\rho_{V}(g_{2})$. Similarly, the Lie algebra representation is defined as $\mathrm{d}\rho_{V}:\mathfrak{g}\rightarrow\mathfrak{gl}(m)$, and for $A_{1},A_{2}\in\mathfrak{g}$, we have $\mathrm{d}\rho_{V}(A_{1}+A_{2})=\mathrm{d}\rho_{V}(A_{1})+\mathrm{d}\rho_{V}(A_{2})$. We can relate the Lie group representation to the Lie algebra representation through the exponential map. Specifically, for $A\in\mathfrak{g}$, $\rho_{V}(\exp(A))=\exp(\mathrm{d}\rho_{V}(A))$ holds. Then, combining with Eq. 1, the matrix group representation can be written as:
@@ 3 Related work
3.1 Equivariant networks
Equivariant networks have gained significant attention in recent years due to their ability to respect and leverage symmetries in data. GCNNs [9] embed discrete group equivariance into traditional CNNs through group convolutions. Steerable CNNs [10] introduce steerable filters, which provide a more flexible and efficient way to achieve equivariance compared with GCNNs. Subsequently, SFCNNs [52] and E(2)-equivariant steerable CNNs [50] extend GCNNs and steerable CNNs to continuous group equivariance, while 3D Steerable CNNs [51] extend these models to 3D volumetric data. On the other hand, some works use partial differential operators (PDOs) to construct equivariant networks [43, 44, 45, 27, 35]. Furthermore, equivariant self-supervised learning [49, 12, 33, 18, 25] has also achieved outstanding results. Based on these theoretical frameworks, equivariant networks are widely applied in various fields, such as mathematics [59], physics [48], biochemistry [2, 53, 22], and others.
3.2 Equivariant Multi-Layer Perceptrons (EMLP)
EMLP [17] embeds matrix group equivariance into MLPs layerwise. Given the input space $U_{i}$ and output space $U_{o}$, the linear weight matrix $W\in U_{o}\otimes U_{i}^{*}$ should satisfy Eq. 5, i.e., $\forall g\in\widetilde{G},v_{i}\in U_{i}:\rho_{o}(g)Wv_{i}=W\rho_{i}(g)v_{i}$. So the coefficients of each term in $v_{i}$ are equal $\forall g\in\widetilde{G}:\rho_{o}(g)W=W\rho_{i}(g)$. Flattening the linear weight matrix $W$ into a vector, we have $\forall g\in\widetilde{G}:\left[\rho_{o}(g)\otimes\rho_{i}(g^{-1})^{\top}\right]\mathrm{vec}(W)=\mathrm{vec}(W)$. Combined with Eq. 4, the linear weight matrix $W$ is invariant in the space $U_{o}\otimes U_{i}^{*}$:
where $\rho_{o,i}=\rho_{o}\otimes\rho_{i}^{*}$ is the group representation of $U_{o}\otimes U_{i}^{*}$. Decomposing the group representation $\rho_{o,i}(g)$ into discrete and infinitesimal generators as shown in Eq. 2, Eq. 6 is equivalent to the following constraint:
3.3 Kolmogorov-Arnold Networks (KANs)
KANs [37, 36] place learnable activation functions on the edges and then sum them to obtain the output nodes, replacing the fixed activation functions applied to the output nodes of linear layers in MLPs. Formally, the $l$-th KAN layer can be expressed as:
@@ 3.3 Kolmogorov-Arnold Networks (KANs)
where $n_{l}$ is the number of nodes in the $l$-th layer, $x_{l,i}$ is the value of the $i$-th node in the $l$-th layer, and $\phi_{l,j,i}$ is the activation function that connects $x_{l,i}$ to $x_{l+1,j}$. In practice, $\phi_{l,j,i}$ consists of a spline function and a $\mathrm{silu}$ function. The spline basis functions are determined by grids, which are updated based on the input samples. Then we can write the post-activation of $\phi_{l,j,i}$ as $\phi_{l,j,i}(x_{l,i})=\sum_{b=0}^{G+k-1}w_{l,j,i,b}B_{l,i,b}(x_{l,i})+w_{l,j,i,G+k}\mathrm{silu}(x_{l,i})$, where $G$, $k$, and $B_{l,i,b}$ represent the number of grid intervals, the order, and the $b$-th basis function of splines at node $x_{l,j}$, respectively. Therefore, the KAN layer can be viewed as spline basis functions $B_{l,i,b}$ and a $\mathrm{silu}$ function, followed by a linear layer with $w_{l,j,i,b}$ as parameters [16]:
@@ 4 EKAN layer
In this section, we construct the EKAN layer, which is equivariant with respect to the matrix group $\widetilde{G}$. First, we define the space structures and explain their relationships. Then, we introduce gated basis functions and equivariant linear weights, which together form a layer of EKAN. Finally, we present the implementation details of grid updates. We summarize the space structures and network architecture of the EKAN layer in Fig. 2.
@@ 4.1 Space structures
We specify the input space and the output space of EKAN layer as $U_{i}$ and $U_{o}$, respectively. Their structures can be normalized into the form of Eq. 3. In particular, for ease of later discussion, we extract the scalar space terms $T_{0}=T(0,0)$ and rewrite them as:
@@
We have to emphasize that the actual input/output feature does not lie within $U_{i}$/$U_{o}$ (the symbol ‘///’ here represents ‘or’ rather than the quotient space). To align with gated basis functions, we add a gate scalar $T_{0}$ to each non-scalar term $T(p_{i,a},q_{i,a})$/$T(p_{o,a},q_{o,a})$ in $U_{i}$/$U_{o}$ to obtain the actual input/output space. The mechanism behind this approach will be discussed in detail in Sec. 4.2. We denote this actual input/output space as the gated input/output space $U_{gi}$/$U_{go}$:
@@
As shown in Sec. 3.3, we split the KAN layer into basis functions and linear weights. From this perspective, we correspondingly construct gated basis functions and equivariant linear weights to form the EKAN layer. We refer to the space where the activation values reside after gated basis functions and before equivariant linear weights as the post-activation space $U_{m}$. The structure of $U_{m}$ depends on $U_{i}$, which we will elaborate on in Sec. 4.2.
We summarize the aforementioned space structures and their relationships in Fig. 2 (Left). The user first specifies the input space $U_{i}$ and the output space $U_{o}$ for the EKAN layer. Then the gated input space $U_{gi}$ and the post-activation space $U_{m}$ are calculated based on $U_{i}$, and the gated output space $U_{go}$ is calculated based on $U_{o}$. The actual input feature in $U_{gi}$ passes through gated basis functions to obtain the activation value in $U_{m}$, which then passes through equivariant linear weights to obtain the actual output feature in $U_{go}$.
@@ 4.2 Gated basis functions
Since spline basis functions are nonlinear and have relatively complex iterative forms, directly solving the equivariance constraint in Eq. 5 is quite challenging. Recently, gating mechanisms have been widely used in various areas, not only in language models to improve performance [13, 42, 24] but also in the design of equivariant networks [51, 17]. Inspired by this, we propose gated basis functions to make the basis functions in KANs (spline basis functions and the $\mathrm{silu}$ function) equivariant. Suppose that the input feature $v_{gi}\in U_{gi}$ can be decomposed according to the space structure shown in Eq. 11:
@@
The following theorem guarantees the equivariance between the gated input space and the post-activation space (see Supplementary Material for proof).
Theorem 1.
Given a matrix group $\widetilde{G}$, the gated input space $U_{gi}$ and the post-activation space $U_{m}$ can be expressed in the forms of Eqs. 11 and 15, respectively. The function $f:U_{gi}\rightarrow U_{m}$ is defined by Eqs. 12, 13 and 14, that is, $v_{m}=f(v_{gi})$. Then, $f$ is equivariant:

### v3

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have seen great success in scientific domains thanks to spline activation functions, becoming an alternative to Multi-Layer Perceptrons (MLPs). However, spline functions may not respect symmetry in tasks, which is crucial prior knowledge in machine learning. In this paper, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), a method for incorporating arbitrary matrix group equivariance into KANs, aiming to broaden their applicability to more fields. We first construct gated spline basis functions, which form the EKAN layer together with equivariant linear weights, and then define a lift layer to align the input space of EKAN with the feature space of the dataset, thereby building the entire EKAN architecture. Compared with baseline models, EKAN achieves higher accuracy with smaller datasets or fewer parameters on symmetry-related tasks, such as particle scattering and the three-body problem, often reducing test MSE by several orders of magnitude. Even in non-symbolic formula scenarios, such as top quark tagging with three jet constituents, EKAN achieves comparable results with state-of-the-art equivariant architectures using fewer than $40\%$ of the parameters, while KANs do not outperform MLPs as expected.
Equivariant neural network, Kolmogorov-Arnold network, Lie theory
@@ 1 Introduction
Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024b, a) are a novel type of neural network inspired by the Kolmogorov-Arnold representation theorem (Tikhomirov, 1991; Braun & Griebel, 2009), which offers an alternative to Multi-Layer Perceptrons (MLPs) (Haykin, 1998; Cybenko, 1989; Hornik et al., 1989). Unlike MLPs, which utilize fixed activation functions on nodes, KANs employ learnable activation functions on edges, replacing the linear weight parameters entirely with univariate functions parameterized as splines (De Boor & De Boor, 1978). On the other hand, each layer of KANs can be viewed as spline basis functions followed by a linear layer (Dhiman, 2024). This architecture allows KANs to achieve better accuracy in symbolic formula representation tasks compared with MLPs, particularly in function fitting and scientific applications. Subsequent works based on KANs have demonstrated superior performance in other areas, such as sequential data (Vaca-Rubio et al., 2024; Genet & Inzirillo, 2024b, a; Xu et al., 2024), graph data (Bresson et al., 2024; De Carlo et al., 2024; Kiamari et al., 2024; Zhang & Zhang, 2024), image data (Cheon, 2024b, a; Azam & Akhtar, 2024; Li et al., 2024a; Seydi, 2024; Bodner et al., 2024), and so on.
@@
However, KANs themselves perform poorly on non-symbolic formula representation tasks (Yu et al., 2024). One reason for this is that splines struggle to respect data type and symmetry, both of which play important roles in machine learning. Many recent works utilize symmetry in data to design network architectures, achieving better efficiency and generalization on specific tasks. For example, Convolutional Neural Networks (CNNs) (LeCun et al., 1989) and Group equivariant Convolutional Neural Networks (GCNNs) (Cohen & Welling, 2016a) leverage translational and rotational symmetries in image data, while DeepSets (Zaheer et al., 2017) and equivariant graph networks (Maron et al., 2018; Keriven & Peyré, 2019; Satorras et al., 2021) exploit the permutation symmetry in set and graph data. Equivariant Multi-Layer Perceptrons (EMLP) (Finzi et al., 2021) propose a general method that allows MLPs to be equivariant with respect to arbitrary matrix groups for specific data types, thereby unifying the aforementioned specialized network architectures.
@@
Inspired by these equivariant architectures, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), which embed matrix group equivariance into KANs. By specifying the data type and symmetry, EKAN can serve as a general framework for applying KANs to various areas. In Section 2, we introduce the preliminary knowledge of group theory. In Section 3, we summarize related works. In Section 4, we construct a layer of EKAN. We add gate scalars to the input and output space of each layer, and define gated spline basis functions between the input and post-activation space. To ensure equivariance when linearly combining gated basis functions, we construct the equivariant constraint and solve for the equivariant linear weights. In Section 5, we build the entire EKAN architecture. We insert a lift layer before the first layer and discard the gate scalars from the output of the final layer, so that the input and output space of EKAN can be consistent with the original dataset. In Section 6, we evaluate EKAN on tasks with known symmetries. We show that EKAN can achieve higher accuracy than baseline models with smaller datasets or fewer parameters. In Section 7, we conclude this work. In Figure 1, we compare the architectures of KANs and EKAN.
@@
Figure 1: Comparison of the architectures of Kolmogorov-Arnold Networks (KANs) and Equivariant Kolmogorov-Arnold Networks (EKAN).
@@
In summary, our contributions are as follows: (1) We propose EKAN, an architecture that makes KANs equivariant to matrix groups. To our knowledge, EKAN is the first attempt to combine equivariance with KANs, and we expect that it can serve as a general framework to broaden the applicability of KANs to more areas. (2) We specify the space structures of the EKAN Layer and define gated spline basis functions. We theoretically prove that gated basis functions can ensure equivariance between the gated input space and the post-activation space. Then, we insert a lift layer to preprocess the raw input feature, which aligns the input space of EKAN with the feature space of the dataset. (3) Experiments on tasks with matrix group equivariance, such as particle scattering and the three-body problem, demonstrate that EKAN often significantly outperforms baseline models, even with smaller datasets or fewer parameters. In the task of non-symbolic formula representation, where KANs are not proficient, such as top quark tagging with three jet constituents, EKAN can still achieve comparable results with state-of-the-art equivariant architectures while using fewer than $40\%$ of the parameters.
@@ 2 Background
where $\{A_{i}\}_{i=1}^{D}$ are called infinitesimal generators and $\{h_{i}\}_{i=1}^{M}$ are called discrete generators. We introduce common matrix groups and their generators in Appendix A.
Group representations.
The group representation $\rho_{V}:\widetilde{G}\rightarrow\mathrm{GL}(m)$ maps group elements to $m\times m$ invertible matrices, which describes how group elements act on the vector space $V=\mathbb{R}^{m}$ through linear transformations. For $g_{1},g_{2}\in\widetilde{G}$ it satisfies $\rho_{V}(g_{1}g_{2})=\rho_{V}(g_{1})\rho_{V}(g_{2})$. Similarly, the Lie algebra representation is defined as $\mathrm{d}\rho_{V}:\mathfrak{g}\rightarrow\mathfrak{gl}(m)$, and for $A_{1},A_{2}\in\mathfrak{g}$, we have $\mathrm{d}\rho_{V}(A_{1}+A_{2})=\mathrm{d}\rho_{V}(A_{1})+\mathrm{d}\rho_{V}(A_{2})$. We can relate the Lie group representation to the Lie algebra representation through the exponential map. Specifically, for $A\in\mathfrak{g}$, $\rho_{V}(\exp(A))=\exp(\mathrm{d}\rho_{V}(A))$ holds. Then, combining with Equation 1, the matrix group representation can be written as:
@@ 3 Related work
Equivariant networks.
Equivariant networks have gained significant attention in recent years due to their ability to respect and leverage symmetries in data. GCNNs (Cohen & Welling, 2016a) embed discrete group equivariance into traditional CNNs through group convolutions. Steerable CNNs (Cohen & Welling, 2016b) introduce steerable filters, which provide a more flexible and efficient way to achieve equivariance compared with GCNNs. Subsequently, SFCNNs (Weiler et al., 2018b) and E(2)-equivariant steerable CNNs (Weiler & Cesa, 2019) extend GCNNs and steerable CNNs to continuous group equivariance, while 3D Steerable CNNs (Weiler et al., 2018a) extend these models to 3D volumetric data. On the other hand, some works use partial differential operators (PDOs) to construct equivariant networks (Shen et al., 2020, 2021, 2022; He et al., 2022; Li et al., 2024b). Furthermore, equivariant self-supervised learning (Wang et al., 2020b; Dangovski et al., 2021; Lee et al., 2022; Garrido et al., 2023; Gupta et al., 2023) has also achieved outstanding results. Based on these theoretical frameworks, equivariant networks are widely applied in various fields, such as mathematics (Zhao et al., 2022), physics (Wang et al., 2020a), biochemistry (Bekkers et al., 2018; Winkels & Cohen, 2019; Graham et al., 2020), and others. EMLP (Finzi et al., 2021) embeds matrix group equivariance into MLPs layerwise, which we discuss in detail in Appendix B.
Kolmogorov-Arnold Networks (KANs).
KANs (Liu et al., 2024b, a) place learnable activation functions on the edges and then sum them to obtain the output nodes, replacing the fixed activation functions applied to the output nodes of linear layers in MLPs. Formally, the $l$-th KAN layer can be expressed as:
@@
where $n_{l}$ is the number of nodes in the $l$-th layer, $x_{l,i}$ is the value of the $i$-th node in the $l$-th layer, and $\phi_{l,j,i}$ is the activation function that connects $x_{l,i}$ to $x_{l+1,j}$. In practice, $\phi_{l,j,i}$ consists of a spline function and a $\mathrm{silu}$ function. The spline basis functions are determined by grids, which are updated based on the input samples. Then we can write the post-activation of $\phi_{l,j,i}$ as $\phi_{l,j,i}(x_{l,i})=\sum_{b=0}^{G+k-1}w_{l,j,i,b}B_{l,i,b}(x_{l,i})+w_{l,j,i,G+k}\mathrm{silu}(x_{l,i})$, where $G$, $k$, and $B_{l,i,b}$ represent the number of grid intervals, the order, and the $b$-th basis function of splines at node $x_{l,j}$, respectively. Therefore, the KAN layer can be viewed as spline basis functions $B_{l,i,b}$ and a $\mathrm{silu}$ function, followed by a linear layer with $w_{l,j,i,b}$ as parameters (Dhiman, 2024):
@@ 4 EKAN layer
In this section, we construct the EKAN layer, which is equivariant with respect to the matrix group $\widetilde{G}$. First, we define the space structures and explain their relationships. Then, we introduce gated basis functions and equivariant linear weights, which together form a layer of EKAN. We summarize the space structures and network architecture of the EKAN layer in Figure 2.
@@ 4.1 Space structures
We specify the input space and the output space of EKAN layer as $U_{i}$ and $U_{o}$, respectively. Their structures can be normalized into the form of Equation 3. In particular, for ease of later discussion, we extract the scalar space terms $T_{0}=T(0,0)$ and rewrite them as:
@@
We have to emphasize that the actual input/output feature does not lie within $U_{i}$/$U_{o}$. To align with gated basis functions, we add a gate scalar $T_{0}$ to each non-scalar term $T(p_{i,a},q_{i,a})$/$T(p_{o,a},q_{o,a})$ in $U_{i}$/$U_{o}$ to obtain the actual input/output space. The mechanism behind this approach will be discussed in detail in Section 4.2. We denote this actual input/output space as the gated input/output space $U_{gi}$/$U_{go}$:
@@
As shown in Section 3, we split the KAN layer into basis functions and linear weights. From this perspective, we correspondingly construct gated basis functions and equivariant linear weights to form the EKAN layer. We refer to the space where the activation values reside after gated basis functions and before equivariant linear weights as the post-activation space $U_{m}$. The structure of $U_{m}$ depends on $U_{i}$, which we will elaborate on in Section 4.2.
We summarize the aforementioned space structures and their relationships in Figure 2 (Left). The user first specifies the input space $U_{i}$ and the output space $U_{o}$ for the EKAN layer. Then the gated input space $U_{gi}$ and the post-activation space $U_{m}$ are calculated based on $U_{i}$, and the gated output space $U_{go}$ is calculated based on $U_{o}$. The actual input feature in $U_{gi}$ passes through gated basis functions to obtain the activation value in $U_{m}$, which then passes through equivariant linear weights to obtain the actual output feature in $U_{go}$.
@@ 4.2 Gated basis functions
Since spline basis functions are nonlinear and have relatively complex iterative forms, directly solving the equivariance constraint in Equation 5 is quite challenging. Recently, gating mechanisms have been widely used in various areas, not only in language models to improve performance (Dauphin et al., 2017; Shazeer, 2020; Gu & Dao, 2023) but also in the design of equivariant networks (Weiler et al., 2018a; Finzi et al., 2021). Inspired by this, we propose gated basis functions to make the basis functions in KANs (spline basis functions and the $\mathrm{silu}$ function) equivariant. Suppose that the input feature $v_{gi}\in U_{gi}$ can be decomposed according to the space structure shown in Equation 9:
@@
The following theorem guarantees the equivariance between the gated input space and the post-activation space (see Appendix C for proof).
Theorem 4.1.
Given a matrix group $\widetilde{G}$, the gated input space $U_{gi}$ and the post-activation space $U_{m}$ can be expressed in the forms of Equations 9 and 13, respectively. The function $f:U_{gi}\rightarrow U_{m}$ is defined by Equations 10, 11 and 12, that is, $v_{m}=f(v_{gi})$. Then, $f$ is equivariant: